Two 2.3-cm diameter disks with charge ±25.0nCface each other, 1.0mm apart.

a. What is the magnitude of the electric field between the disks?

b. A proton is shot from the negative disk toward the positive disk. What launch speed must it have to just barely reach the positive disk?
Answer #1

A)

Electric field, E = Q/(A e0)

E = (25 x 10^-9)/(8.85 x 10^-12 x pi (0.023/2)^2)

E = 6.80 x 10^6 N/C

B)

Acceleration, a = qE/m

a = (1.6 x 10^-19 x 6.8 x 10^6)/(9.1 x 10^-31) = 1.20 x 10^18 m/s^2

v^2 = u^2 + 2as

v^2 = 0 + 2 x 1.2 x 10^18 x 0.001

V = 4.89 x 10^7 m/s

Comment in case any doubt please rate my answer....